# Over-Squashing in Graph Neural Networks: A Comprehensive survey

### v6

@@ V Graph Transformers and other Strategies
TABLE II: Dataset Statistics
@@
Cora Citeseer Pubmed Film TwitchDE Tolokers Cornell Texas Wisconsin Chemeleon Squirrel Actor
#nodes 2708 3327 19717 7600 9496 11758 183 183 251 2277 5201 7600
#edges 10556 9104 88648 53504 153138 519000 280 295 466 31421 198493 26752
#features 1433 3703 500 932 2514 10 1703 1703 1703 2089 2089 932
#classes 7 6 3 5 2 2 5 5 5 5 5 5
$H(G)$ 0.81 0.74 0.80 0.22 0.63 0.59 0.30 0.11 0.21 0.23 0.22 0.22
@@ VI Datasets
The common datasets employed for node and graph classification tasks in the models listed in Table I are presented in Table II, along with detailed dataset statistics. It’s important to note that this list is not exhaustive, as there are numerous other datasets, including synthetic and large-scale real-world ones, utilized for various research purposes. Table II displays the statistics of the datasets used in this study, where $H(G)$ represents the graph’s homophily, as defined in [84], calculated as
@@
Cora [85], CiteSeer [86], and PubMed [87] are examples of paper citation networks. In these datasets, node features are represented as bag-of-words extracted from paper content, and the goal is to classify research topics. Notably, these datasets exhibit high homophily. In contrast, the Film dataset is created based on actor co-occurrences on Wikipedia pages and is categorized into five groups. This dataset poses a node classification task with low homophily characteristics. TwitchDE, on the other hand, is a social network comprising German gamer accounts from Twitch, categorized as suitable for work or adult profiles. The classification task involves profiling these accounts. The Tolokers dataset represents a collaboration network derived from the crowdsourcing platform Toloka. The objective here is to determine user activity, considering the challenge of class imbalance, with the evaluation metric being the area under the ROC curve. Cornell, Texas, Wisconsin are additional node classification tasks, each originating from university-related interactions. The Cornell dataset comprises research paper citation data. The Texas dataset represents friendships in a Texas college, and the Wisconsin dataset is derived from a university-related network. Node features and specific targets for these datasets can vary. Chameleon[88], Squirrel[88], Actor [89] are also novel datasets introduced for node classification. Chameleon captures interactions within a university community. Squirrel is a network of interactions among squirrels in a park. The Actor dataset models collaborations among actors in the film industry. Each of these datasets presents unique characteristics and classification tasks.
@@
In all these tasks, we intentionally avoid introducing structural input features such as node degrees or positional encodings. A summary of relevant dataset statistics is provided in Table II for reference.
@@ VII Conclusion and Future Direction
As discussed in this survey, while techniques exist to mitigate the effects of over-squashing in GNNs, several directions remain open for exploration:
@@
    Dynamic Graph Adaptation: Existing models and rewiring methods in the literature are primarily designed for static graphs, overlooking the dynamic nature of real-world graph data. Investigating the performance of these techniques, as well as developing novel approaches, in the realm of dynamic graphs is essential.
@@
    Handling Distributional Shifts: Graph representation learning often struggles with generalizing to out-of-distribution data, leading to performance degradation in real-world scenarios. While techniques exist to mitigate distributional shifts in graph data, the impact of over-squashing in such contexts remains underexplored.
@@
    Robustness to Noisy Data: Investigating techniques to improve the robustness of GNNs to noisy or adversarial data, which can exacerbate issues such as over-squashing. This may involve exploring methods for robust training or designing inherently resilient models to perturbations in the input data.
@@
    Scalability and Efficiency: Addressing scalability and efficiency concerns associated with existing over-squashing mitigation techniques, especially for large-scale graph datasets. This may entail developing scalable algorithms or designing lightweight model architectures suitable for deployment in resource-constrained environments.
  5. 5.
    Incorporating Heterophily in Graph Classification: While discussed models incorporate heterophily only in node classification tasks, exploring its impact on graph classification tasks would be intriguing. Analyzing the role of heterophily in graph classification could lead to insights into improving the performance and robustness of graph classification models.

### v7

@@ IV Unifying Approaches to Address Over-Squashing and Over-Smoothing Trade-offs in Graph Neural Networks
Several methods have been proposed to address over-smoothing and over-squashing in GNNs by aggregating features from multiple neighborhood hops using weighted combinations of adjacency matrix powers. Models such as SGC [87], S2GC [88], and APPNP [89] follow this approach, while GDC [51] generalizes the graph diffusion process to provide a unified framework. These methods extend GCN depth without severe performance degradation; however, they inevitably sacrifice some long-range information to alleviate over-smoothing, resulting in a tradeoff between mitigating over-smoothing and addressing over-squashing [32].
To overcome this limitation, Shen et al. [70] proposed the Maximization-based Graph Convolution (MGC) model. Unlike traditional models that rely on linear aggregation of multi-hop neighborhood features, MGC employs an elementwise maximization operation to gather information from all powers of the adjacency matrix. This approach improves information flow across distant nodes, effectively addressing both over-smoothing and over-squashing. To improve scalability for large graphs, the authors introduced an approximated variant (AMGC) with linear complexity in the number of edges. MGC follows a decoupled design where feature propagation occurs in a preprocessing stage, followed by a simple MLP-based transformation during training. By combining original and processed node features, this architecture enhances model expressiveness while maintaining computational efficiency.
While MGC focuses on improving aggregation strategies, Mancini and Rekik [90] address over-smoothing and over-squashing through an edge-filtering approach. Their proposed model, DuoGNN, introduces a scalable and generalizable GNN architecture that decouples homophilic and heterophilic interactions in a topological interaction-decoupling stage. This separation ensures that each type of interaction is processed independently. DuoGNN follows three key steps: First, a topological edge-filtering step removes heterophilic edges to construct a strongly homophilic subgraph $G_{ho}$, preventing over-smoothing and preserving short-range interactions. Second, a heterophilic graph condensation step extracts important heterophilic interactions from homophilic clusters to form a new subgraph $G_{he}$ which captures long-range dependencies while maintaining a compact structure. Finally, DuoGNN employs two independent GNN modules that process $G_{ho}$ and $G_{he}$ separately—one learning homophilic aggregation and the other modeling heterophilic interactions. This dual-path design enables DuoGNN to balance expressiveness and scalability effectively.
Building upon concepts from DuoGNN, DeltaGNN [91] extends this idea by introducing a more structured Interaction Filtering and Clustering (IFC) mechanism. Similar to DuoGNN, DeltaGNN begins by performing homophilic aggregation to capture short-range dependencies. However, rather than relying on fixed connectivity rules, DeltaGNN dynamically removes edges with the lowest IFC scores, calculated using Euclidean distance. This targeted edge removal refines the graph structure by eliminating bottlenecks while preserving homophilic components. To restore long-range dependencies lost during this filtering, DeltaGNN employs a heterophilic graph condensation process similar to DuoGNN’s strategy. However, DeltaGNN identifies key nodes with high IFC scores to build this graph, ensuring efficient retention of critical long-range interactions. Both DuoGNN and DeltaGNN adopt a dual aggregation strategy, processing homophilic and heterophilic graphs independently. While DuoGNN emphasizes graph topology to split interactions, DeltaGNN’s adaptive filtering process enhances scalability and performance on complex datasets.
@@ V Graph Transformers and other Strategies
TABLE V: Dataset Statistics
@@
Cora Citeseer Pubmed Actor TwitchDE Tolokers Cornell Texas Wisconsin Chemeleon Squirrel Roman-Empire
#nodes 2708 3327 19717 7600 9496 11758 183 183 251 2277 5201 22662
#edges 10556 9104 88648 26752 153138 519000 280 295 466 31421 198493 32927
#features 1433 3703 500 932 2514 10 1703 1703 1703 2089 2089 300
#classes 7 6 3 5 2 2 5 5 5 5 5 18
$H(G)$ 0.81 0.74 0.80 0.22 0.63 0.59 0.30 0.11 0.21 0.23 0.22 0.06
@@ VI Datasets
The common datasets employed for node and graph classification tasks in the models listed in Table I are presented in Table V, along with detailed dataset statistics. It’s important to note that this list is not exhaustive, as there are numerous other datasets, including synthetic and large-scale real-world ones, utilized for various research purposes. Table V displays the statistics of the datasets used in this study, where $H(G)$ represents the graph’s homophily, as defined in [99], calculated as
@@
Cora [100], CiteSeer [101], and PubMed [102] are examples of paper citation networks. In these datasets, node features are represented as bag-of-words extracted from paper content, and the goal is to classify research topics. Notably, these datasets exhibit high homophily. In contrast, the Actor [103] dataset is created based on actor co-occurrences on Wikipedia pages and is categorized into five groups. This dataset poses a node classification task with low homophily characteristics. TwitchDE, on the other hand, is a social network comprising German gamer accounts from Twitch, categorized as suitable for work or adult profiles. The classification task involves profiling these accounts. The Tolokers dataset represents a collaboration network derived from the crowdsourcing platform Toloka. The objective here is to determine user activity, considering the challenge of class imbalance, with the evaluation metric being the area under the ROC curve. Cornell, Texas, Wisconsin are additional node classification tasks, each originating from university-related interactions. The Cornell dataset comprises research paper citation data. The Texas dataset represents friendships in a Texas college, and the Wisconsin dataset is derived from a university-related network. Node features and specific targets for these datasets can vary. Chameleon[104], Squirrel[104] and Roman-Empire [105] are also novel datasets introduced for node classification. Chameleon captures interactions within a university community. Squirrel is a network of interactions among squirrels in a park. The Roman-Empire dataset represents a word co-occurrence network derived from the Roman Empire Wikipedia article. In this dataset, nodes correspond to individual words, and edges are formed between words that are either sequentially linked or syntactically related. Since the edges follow a specific order reflecting the sequence or syntactic structure of the text, the Roman-Empire dataset is inherently a directed graph.
@@
In all these tasks, we intentionally avoid introducing structural input features such as node degrees or positional encodings. A summary of relevant dataset statistics is provided in Table V for reference.
@@ VII Conclusion and Future Direction
While this survey highlights various techniques to mitigate the effects of over-squashing in GNNs, several promising directions remain open for further exploration:
@@
    Dynamic Graph Adaptation: Most existing models and rewiring techniques are designed for static graphs, where the node and edge structures remain fixed throughout the learning process. Although effective in certain scenarios, these methods often overlook the dynamic nature of real-world graphs. In static graphs, rewiring methods mitigate over-squashing by strategically adding edges or modifying the graph structure to improve long-range information flow. However, these fixed modifications may prove insufficient for dynamic graphs, where evolving connectivity patterns require adaptive solutions. To address this, Petrovi’c et al. [107] proposed a Temporal Graph Rewiring (TGR) method specifically for dynamic graphs. Unlike static rewiring, TGR dynamically refines connectivity patterns by introducing time-aware edges that optimize information flow as the graph evolves. This adaptive strategy effectively reduces the risk of over-squashing, even in rapidly changing graph structures. While dynamic graphs offer flexibility in maintaining efficient information propagation, they also introduce practical challenges. Tracking and updating structural changes in real-time demands additional computational resources, making scalability a critical concern for dynamic graph rewiring strategies.
@@
    Handling Distributional Shifts: Graph representation learning often struggles to generalize effectively to out-of-distribution data, resulting in performance degradation in real-world scenarios. While certain techniques address distributional shifts in graph data, the interplay between these shifts and the over-squashing problem remains underexplored. Future research could investigate how mitigating over-squashing might enhance model robustness under such conditions.
@@
    Robustness to Noisy Data: Noise and adversarial perturbations in graph data can amplify the effects of over-squashing, further impairing model performance. Exploring techniques that improve GNN robustness — such as robust training methods or models inherently resilient to data perturbations — is crucial for enhancing performance in noisy environments.
@@
    Scalability and Efficiency: Existing over-squashing mitigation techniques often introduce additional computational overhead, posing challenges for large-scale graph datasets. Developing scalable algorithms or lightweight architectures that effectively mitigate over-squashing without excessive resource demands remains an important research direction.